OpenStudy (aravindg):

let f(x)=x^2 and g(x)=sinx for all x belongs to R show that gof not equal to fog

OpenStudy (aravindg):

giv answer plzz

OpenStudy (anonymous):

what course are u taking?

OpenStudy (aravindg):

???

OpenStudy (anonymous):

1. Define f o g and g o f 2. Find one example where f o g (x) != g o f (x)

OpenStudy (anonymous):

isnt this simple substitution

OpenStudy (anonymous):

and comparison

OpenStudy (anonymous):

f o g (x) = sin^2 x g o f (x) = sin x^2 try x = 1 one for example. f o g (1) = sin^2 1 = .708... g o f (1) = sin 1^2 = sin 1 = .841... Proven.

OpenStudy (anonymous):

^thnkx